# Hot, Photoionized X-ray Gas in Two Luminous Type 2 Quasars: Chandra-HST Evidence for a Wind-Driven Sequence

### v1

@@ Abstract
We present new Chandra/ACIS-S imaging spectroscopy of two luminous type 2 quasars, FIRST J120041.4+314745 (=0.116) and 2MASX J13003807+5454367 (=0.088), and compare their X-ray emission with Hubble Space Telescope [O III]5007 morphologies and kinematics. Both systems show kiloparsec-scale soft X-ray emission. In FIRST J120041, the X-ray morphology is clumpy and closely follows the [O III] structures, with surface-brightness peaks co-spatial with the highest [O III] velocities (600–750 km s-1) and broadest line widths (FWHM1700 km s-1). In 2MASX J130038, the X-rays are more centrally concentrated and only weakly correlated with the largely rotational [O III] kinematics. Spectral modeling indicates that photoionization dominates the soft X-rays in both quasars. The inferred hot-gas reservoirs are substantial, (FIRST J120041) and (2MASX J130038), exceeding the outflowing [O III] masses (volume-normalized) by factors of 4 and 16, respectively. In 2MASX J130038, we also identify a tentative blueshifted Fe XXVI Ly line at keV ( km s-1), consistent with a nascent hot wind confined to the inner few hundred parsecs. Combining these results with a broader sample of twelve type 2 quasars, we argue that luminous quasars evolve along a continuous feedback sequence regulated by the progressive clearing of circumnuclear gas. As AGN radiation and winds pierce through the surrounding medium, systems transition from heavily enshrouded, compact configurations to phases where the X-ray and [O III] components strongly couple and, eventually, to large-scale, energetically dominant outflows. FIRST J120041 and 2MASX J130038 represent two points along this sequence, tracing the emergence and growth of hot winds as primary drivers of quasar-scale feedback.
1 Introduction
@@
Active galactic nuclei (AGNs) have the potential to influence the evolution of galaxies through mass outflows that interact with the surrounding interstellar medium (ISM). These feedback processes may regulate the co-evolution of supermassive black holes (SMBHs) and their host galaxies, helping to establish observed scaling relations such as the - correlation between SMBH mass and bulge velocity dispersion (e.g., gebhardt_relationship_2000; begelman_agn_2004). Despite their potential significance, the physical mechanisms that might drive AGN feedback remain under debate, largely because they are difficult to isolate observationally. The relevant processes operate on different spatial and temporal scales, and many potential mechanisms, such as induction heating, shocks, or outflows (e.g., gaspari_unifying_2017; gaspari_linking_2020; trindade_falcao_hubble_2021; falcone_analysis_2024), leave ambiguous or indirect signatures that make it challenging to identify the dominant driver in any given system.
@@
At low luminosities (bolometric luminosities ), nearby () Seyfert galaxies provide valuable insight through narrow-line region (NLR) outflows. Optical spectroscopy of the [O III] line has been central to mapping NLR kinematics, revealing signatures of AGN-driven winds launched from near the SMBH as well as outflows arising from in situ gas entrained within the host galaxy disk (fischer_determining_2013; crenshaw_feedback_2015; fischer_gemini_2017). Radiative acceleration is often invoked with considerable success to explain these outflows (meena_radiative_2021; meena_investigating_2023; falcone_analysis_2024; tutterow_shape_2025), yet the dynamics and origin of NLR winds are still not fully understood (e.g., gaspari_unifying_2017; gaspari_linking_2020).
@@
At higher luminosities, type 2 quasars (QSO2s) provide a unique window into feedback operating on kiloparsec scales. Using Hubble Space Telescope (HST) [O III] imaging and Space Telescope Imaging Spectrograph (STIS) spectroscopy, fischer_hubble_2018 showed that local, luminous QSO2s (, ) can host NLRs extending over several kiloparsecs (see also e.g., storchi-bergmann_bipolar_2018), with clear outflow signatures in their inner regions (see also e.g., de_oliveira_gauging_2021). However, they also found that not all extended [O III] NLRs are accompanied by equally extended [O III] outflows. This discrepancy raises a key question: if outflows are radiatively accelerated and the AGN is luminous enough to ionize gas across kiloparsec scales, why do some outflows appear to stall near the nucleus?
@@
One possibility is that hot X-ray winds provide the dominant dynamical driver, entraining and accelerating cooler gas phases such as the [O III]-emitting component (trindade_falcao_hubble_2021). Indeed, ultra-fast outflows (UFOs) detected in X-rays have been linked to galaxy-scale molecular winds in ultra-luminous infrared galaxies (ULIRGs; tombesi_wind_2015), suggesting a physical connection between hot winds and cooler outflowing gas. In nearby Seyferts, Chandra imaging reveal extended soft X-ray emission that is spatially coincident with [O III] structures (e.g., young_chandra_2001; wang_deep_2011; maksym_cheers_2019), while blueshifted X-ray emission lines show velocities comparable to those observed in the optical gas (e.g., kraemer_physical_2015; kallman_census_2014). In some systems, the X-ray outflow extends even beyond the optical NLR, implying that the hot phase may dominate the total energy budget (kraemer_mass_2020).
@@
These correlations persist at higher luminosities. The luminous QSO2 Mrk 34, for example, shows a well-defined [O III] bicone and high-velocity outflows () extending to pc (fischer_hubble_2018). Multi-component modeling of its NLR (revalski_quantifying_2018; revalski_erratum_2019) combined with spatially resolved STIS spectroscopy yielded a peak outflow rate at kpc and peak kinetic luminosity at pc (trindade_falcao_hubble_2021-1). The close morphological agreement between [O III] and soft X-ray emission in Mrk 34 reinforces the idea of a physical coupling between these phases.
@@
To investigate this connection further, we obtained new Chandra/ACIS-S observations of two luminous QSO2s from the fischer_hubble_2018 sample: FIRST J120041.4+314745 (, 2.04 kpc arcsec-1; hereafter FIRST J120041) and 2MASX J13003807+5454367 (, 1.59 kpc arcsec-1; hereafter 2MASX J130038). These objects are the two [O III]5007-brightest ( and 42.7, respectively; trindade_falcao_hubble_2021-1) of the extended sources following Mrk 34 and have the largest deprojected NLRs ( kpc and 4.7 kpc, respectively; fischer_hubble_2018). Despite their similar NLR sizes, their [O III] outflows show strikingly different kinematics: FIRST J120041 hosts high-velocity outflows extending to 1 kpc, whereas 2MASX J130038 shows predominantly rotational motion. Photoionization modeling of their STIS spectra underscores this contrast, with FIRST J120041 showing [O III] mass outflow rates and kinetic luminosities that exceed those of 2MASX J130038 by factors of 210 and 100, respectively (trindade_falcao_hubble_2021-1).
@@
2 Observations and Data Reduction
@@
We retrieved the raw event files from the Chandra Data Archive111https://cda.harvard.edu/chaser/ and reprocessed them with CIAO 4.17222http://cxc.harvard.edu/ciao (fruscione_ciao_2006), applying the latest calibration files through the chandra_repro333https://cxc.cfa.harvard.edu/ciao/ahelp/chandra_repro.html script. Subpixel event repositioning was enabled to fully exploit the angular resolution of Chandra and resolve extended structures on sub-arcsecond scales.
To ensure precise registration among individual observations, we refined the astrometry using the fine_astro444https://cxc.cfa.harvard.edu/ciao/ahelp/fine_astro.html routine, which cross-matches detected X-ray sources with an external reference catalog and applies small linear shifts to the event files. The resulting positional offsets (dx, dy) in units of sky pixels are listed in Table 1 and are consistent with the nominal uncertainties of the Chandra aspect solution. The reprocessed and aligned event files were then merged to produce the final datasets used for the imaging analysis described in Section 4.
Spectra were extracted using the specextract555https://cxc.cfa.harvard.edu/ciao/ahelp/specextract.html task. For FIRST J120041, we used a circular aperture of radius centered at R.A.=12:00:41.4456, decl.=+31:47:46.062, while for 2MASX J130038 we adopted a aperture centered at R.A.=13:00:38.2106, decl.=+54:54:36.739. These apertures encompass the full extent of the detected emission. Background spectra were extracted from nearby, source-free regions with radii of .
@@
3 Spectral Properties of the X-ray Emission
3.1 Phenomenological Models
We first model the spectra using phenomenological, multi-component fits implemented in Sherpa777https://cxc.cfa.harvard.edu/sherpa/. These models are intended to identify the main continuum components and locate prominent emission features without imposing strong physical assumptions (e.g., trindade_falcao_deep_2023).
The baseline model consists of an absorption component (xstbabs) allowed to vary above the Galactic value888calculated using the HEASARC tool: https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl ( cm-2 for FIRST J120041, and cm-2 for 2MASX J130038), combined with a series of redshifted Gaussian lines (xszgauss) modeling discrete emission features. When required by the data, we include a hard X-ray continuum component 3 keV, modeled either as a simple power law (xspowerlaw) or as a physically motivated reflection component using MYTorus999https://www.mytorus.com/ (murphy_x-ray_2009). In the MYTorus configuration, the scattered continuum (myts) and fluorescent line (mytl) components were linked to a common column density, with the inclination fixed at 90∘. Owing to limited photon statistics, the photon index is fixed at for all MYTorus and xspowerlaw components.
@@ 3.1 Phenomenological Models
The spectrum shows a pronounced hump between 3-6 keV, along with distinct emission features at both lower and higher energies. The best-fit model (Cstat/d.o.f. = 0.76) requires a myts scattered continuum, its associated mytl fluorescent line component, and an additional Gaussian at keV. Below 3 keV, several Gaussian lines are detected at keV (Mg XI, possibly blended with Mg XII), keV (Si XIII/Si XIV), keV (S XV/S XVI), and keV (S XV). The 7.14 keV feature is discussed further in Section 6.2 as a candidate blueshifted Fe XXVI Ly line. The best-fit line-of-sight absorption is cm-2.
3.2 Physically Motivated Models
To investigate the physical origin of the X-ray emission, we next fit models composed of photoionized and, when necessary, thermal plasma components. The photoionized emission is modeled using Cloudy (chatzikos_2023_2023), while thermal emission is tested with Apec (foster_updated_2012). All fits adopt the phenomenological continuum from Section 3.1, with the intrinsic absorbing column allowed to vary above the Galactic value.
The Cloudy grids assume a broken power-law spectral energy distribution (SED) of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5-10 keV, and from 10-100 keV, adopting 1.4 solar abundances (trindade_falcao_hubble_2021-1). The final grids span in steps of 0.25 dex and in 0.1 dex steps. The spectra are exported as XSPEC-style additive tables (porter_cloudyxspec_2006) and fitted using Sherpa.
A physical requirement of the photoionization models is that the emitting slab must be geometrically thin at the distance of the gas. Specifically, the slab thickness should satisfy (e.g., kraemer_physical_2015). This ensures that the cloud does not exceed either the amount of material available at that radius or the maximum emitting surface implied by the morphology. In practice, this condition introduces a coupled constraint between the ionization parameter and the allowed range of column densities : for a given , only certain are physically plausible.
Because varies during the fitting process, the upper limit on the column density, cannot be imposed as a single fixed boundary. Instead, we implement a dynamic constraint in Sherpa that updates the maximum allowed column density at each iteration of the fit. The constraint has the general form
@@ 3.2 Physically Motivated Models
The resulting best-fit physical models are presented in Figure 2, and their parameter values in Table 3.
@@
The spectrum is best described (Cstat/d.o.f. = 1.73) by two photoionized components: one medium-ionization () and medium-column ( cm-2), and one high-ionization () and low-column ( cm-2) component. No additional continuum is required to reproduce the data.
@@
Adding a thermal plasma component does not significantly improve either fit, indicating that photoionization alone adequately accounts for the observed X-ray spectra in both quasars.
@@
4 Morphological Properties of the X-ray Emission
4.1 PSF Modeling
Accurate modeling of the Chandra point spread function (PSF) is essential for distinguishing real extended emission from unresolved nuclear components. For each observation, we simulate the PSF using the Chandra Ray Tracer (ChaRT101010https://cxc.cfa.harvard.edu/ciao/PSFs/chart2/index.html) and projected the rays onto the detector plane with MARX111111https://space.mit.edu/cxc/marx/. We generate ensembles of 1,000 PSF realizations centered at the source position and merge them within the same energy bands adopted for the imaging analysis (Section 4.2).
Following the empirical method of fabbiano_revisiting_2020, each simulated PSF is normalized to the observed counts within an annulus of 0.2′′-0.5′′ radius. This normalization assumes that the observed emission comprises a central point-like component plus any additional extended structure. By testing multiple normalization radii we ensure robust scaling and minimize the risk of over-subtracting flux from the PSF wings.
4.2 Broad- and Narrow-Band Imaging

### v2

@@ Abstract
We present new Chandra/ACIS-S imaging spectroscopy of two luminous type 2 quasars, FIRST J120041.4+314745 (=0.116) and 2MASX J13003807+5454367 (=0.088), and compare their X-ray emission with Hubble Space Telescope [O III]5007 morphologies and kinematics. Both systems show kiloparsec-scale soft X-ray emission. In FIRST J120041, the X-ray morphology is clumpy and closely follows the [O III] structures, with surface-brightness peaks co-spatial with the highest [O III] velocities (600-750 km s-1) and broadest line widths (1700 km s-1). In 2MASX J130038, the X-ray emission is centrally concentrated and weakly correlated with rotational [O III] kinematics. Spectral modeling indicates that photoionization dominates the soft X-rays in both quasars. The inferred hot-gas reservoirs are substantial, (FIRST J120041) and (2MASX J130038), exceeding the outflowing [O III] masses by factors of 4 and 16. In 2MASX J130038, we identify a tentative blueshifted Fe XXVI Ly line at keV ( km s-1), consistent with a hot wind confined to the inner few hundred parsecs. Combining these results with a broader sample of twelve type 2 quasars, we argue that luminous quasars evolve along a continuous feedback sequence regulated by progressive clearing of circumnuclear gas. As AGN radiation and winds pierce the surrounding medium, systems transition from heavily enshrouded, compact configurations to phases where the X-ray and [O III] components strongly couple and, eventually, to energetically dominant outflows. FIRST J120041 and 2MASX J130038 represent two points along this sequence, tracing the emergence and growth of hot winds as primary drivers of quasar-scale feedback.
I Introduction
@@
Active galactic nuclei (AGNs) have the potential to influence the evolution of galaxies through mass outflows that interact with the surrounding interstellar medium (ISM). These feedback processes may regulate the co-evolution of supermassive black holes (SMBHs) and their host galaxies, helping to establish observed scaling relations such as the - correlation between SMBH mass and bulge velocity dispersion (e.g., Gebhardt et al., 2000; Begelman, 2004). Despite their potential significance, the physical mechanisms that might drive AGN feedback remain under debate, largely because they are difficult to isolate observationally. The relevant processes operate on different spatial and temporal scales, and many potential mechanisms, such as induction heating, shocks, or outflows (e.g., Gaspari and Sadowski, 2017; Gaspari et al., 2020; Trindade Falcão et al., 2021a; Falcone et al., 2024), leave ambiguous or indirect signatures that make it challenging to identify the dominant driver in any given system.
@@
At low luminosities (bolometric luminosities ), nearby () Seyfert galaxies provide valuable insight through narrow-line region (NLR) outflows. Optical spectroscopy of the [O III] line has been central to mapping NLR kinematics, revealing signatures of AGN-driven winds launched from near the SMBH as well as outflows arising from in situ gas entrained within the host galaxy disk (Fischer et al., 2013; Crenshaw et al., 2015; Fischer et al., 2017). Radiative acceleration is often invoked with considerable success to explain these outflows (Meena et al., 2021, 2023; Falcone et al., 2024; Tutterow et al., 2025), yet the dynamics and origin of NLR winds are still not fully understood (e.g., Gaspari and Sadowski, 2017; Gaspari et al., 2020).
@@
At higher luminosities, type 2 quasars (QSO2s) provide a unique window into feedback operating on kiloparsec scales. Using Hubble Space Telescope (HST) [O III] imaging and Space Telescope Imaging Spectrograph (STIS) spectroscopy, Fischer et al. (2018) showed that luminous, local QSO2s (, ) can host NLRs extending over several kiloparsecs (see also e.g., Storchi-Bergmann et al., 2018), with clear outflow signatures in their inner regions (see also e.g., de Oliveira et al., 2021). However, they also found that not all extended [O III] NLRs are accompanied by equally extended [O III] outflows. This discrepancy raises a key question: if outflows are radiatively accelerated and the AGN is luminous enough to ionize gas across kiloparsec scales, why do some outflows appear to stall near the nucleus?
@@
One possibility is that hot X-ray winds provide the dominant dynamical driver, entraining and accelerating cooler gas phases such as the [O III]-emitting component (Trindade Falcão et al., 2021a). Indeed, ultra-fast outflows (UFOs) detected in X-rays have been linked to galaxy-scale molecular winds in ultra-luminous infrared galaxies (ULIRGs; Tombesi et al. 2015), suggesting a physical connection between hot winds and cooler outflowing gas. In nearby Seyferts, Chandra imaging reveal extended soft X-ray emission that is spatially coincident with [O III] structures (e.g., Young et al., 2001; Wang et al., 2011; Maksym et al., 2019), while blueshifted X-ray emission lines show velocities comparable to those observed in the optical gas (e.g., Kraemer et al., 2015; Kallman et al., 2014). In some systems, the X-ray outflow extends even beyond the optical NLR, implying that the hot phase may dominate the total energy budget (Kraemer et al., 2020).
@@
These correlations persist at higher luminosities. The luminous QSO2 Mrk 34, for example, shows a well-defined [O III] bicone and high-velocity outflows () extending to pc (Fischer et al., 2018). Multi-component modeling of its NLR (Revalski et al., 2018, 2019) combined with spatially resolved STIS spectroscopy yielded a peak outflow rate at kpc and peak kinetic luminosity at pc (Trindade Falcão et al., 2021b). The close morphological agreement between [O III] and soft X-ray emission in Mrk 34 reinforces the idea of a physical coupling between these phases.
@@
To investigate this connection further, we obtained new Chandra/ACIS-S observations of two luminous QSO2s from the Fischer et al. (2018) sample: FIRST J120041.4+314745 (, 2.04 kpc arcsec-1; hereafter FIRST J120041) and 2MASX J13003807+5454367 (, 1.59 kpc arcsec-1; hereafter 2MASX J130038). These objects are the two [O III]5007-brightest ( and 42.7, respectively; Trindade Falcão et al. 2021b) of the extended sources following Mrk 34 and have the largest deprojected NLRs ( kpc and 4.7 kpc, respectively; Fischer et al. 2018). Despite their similar NLR sizes, their [O III] outflows show strikingly different kinematics: FIRST J120041 hosts high-velocity outflows extending to 1 kpc, whereas 2MASX J130038 shows predominantly rotational motion. Photoionization modeling of their STIS spectra underscores this contrast, with FIRST J120041 showing [O III] mass outflow rates and kinetic luminosities that exceed those of 2MASX J130038 by factors of 210 and 100, respectively (Trindade Falcão et al., 2021b).
@@
II Observations and Data Reduction
@@
We retrieved the raw event files from the Chandra Data Archive111https://cda.harvard.edu/chaser/ and reprocessed them with CIAO 4.17222http://cxc.harvard.edu/ciao (Fruscione et al., 2006), applying the latest calibration files through the chandra_repro333https://cxc.cfa.harvard.edu/ciao/ahelp/chandra_repro.html script. Subpixel event repositioning was enabled to fully exploit the angular resolution of Chandra and resolve extended structures on sub-arcsecond scales.
To ensure precise registration among individual observations, we refined the astrometry using the fine_astro444https://cxc.cfa.harvard.edu/ciao/ahelp/fine_astro.html routine, which cross-matches detected X-ray sources with an external reference catalog and applies small linear shifts to the event files. The resulting positional offsets (dx, dy) in units of sky pixels are listed in Table 1 and are consistent with the nominal uncertainties of the Chandra aspect solution. The reprocessed and aligned event files were then merged to produce the final datasets used for the imaging analysis described in Section IV.
Spectra were extracted using the specextract555https://cxc.cfa.harvard.edu/ciao/ahelp/specextract.html task. For FIRST J120041, we used a circular aperture of radius centered at R.A.=12:00:41.4456, decl.=+31:47:46.062, while for 2MASX J130038 we adopted a aperture centered at R.A.=13:00:38.2106, decl.=+54:54:36.739. These apertures encompass the full extent of the detected emission. Background spectra were extracted from nearby, source-free regions with radii of (centered at R.A.=12:00:43.3043, decl.=+31:48:05.418 for FIRST J120041, and R.A.=13:00:41.4230, decl.=+54:54:53.053 for 2MASX J130038).
@@
III Spectral Properties of the X-ray Emission
III.1 Phenomenological Models
We first model the spectra using phenomenological, multi-component fits implemented in Sherpa777https://cxc.cfa.harvard.edu/sherpa/. These models are intended to identify the main continuum components and locate prominent emission features without imposing strong physical assumptions (e.g., Trindade Falcão et al., 2023).
The baseline model consists of an absorption component (xstbabs) allowed to vary above the Galactic value ( cm-2 for FIRST J120041, and cm-2 for 2MASX J130038, calculated using the HEASARC tool888https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3nh/w3nh.pl, Bekhti et al. 2016), combined with a series of redshifted Gaussian lines (xszgauss) modeling discrete emission features. The Gaussian components have line widths () free to vary, but are required to satisfy 0.1 keV, corresponding to the ACIS-S energy resolution.
When required by the data, we include an additional hard X-ray continuum component at 3 keV to account for nuclear AGN emission. In cases where the absorbed nuclear continuum is sufficiently strong to dominate the hard-band spectrum over the diffuse, non-nuclear emission (as in 2MASX J130038), this component is modeled using the physically motivated torus reflection model MYTorus (Murphy and Yaqoob, 2009). In the MYTorus configuration, the scattered continuum (myts) and fluorescent line (mytl) components are linked to a common column density, with the inclination fixed at 90∘.
In contrast, for sources in which the nuclear emission does not dominate the total X-ray output but is still required to reproduce the observed continuum at 3 keV (as in FIRST J120041), we adopt a simpler absorbed power-law component (xspowerlaw). Owing to limited photon statistics, the photon index is fixed at for all MYTorus and xspowerlaw components.
The phenomenological models were constructed iteratively, beginning with a minimal absorbed continuum and adding components only when statistically and visually warranted by the data. After fitting the baseline model, we introduced Gaussian emission lines one at a time, retaining them only if they produced a significant improvement in the fit statistic and removed coherent residual features.
Because we adopt the Cash statistic (Cstat) for low-count spectra, we evaluate model improvement using the change in C-statistic (Cstat) together with the removal of localized residuals. Gaussian components are retained only when they reduce the C-statistic by Cstat for two additional free parameters and when the associated residuals are repeatable across independent observations.
For nested models, Cstat asymptotically follows a distribution under Wilks’ theorem; for two additional degrees of freedom, Cstat corresponds to an improvement at approximately the 99% confidence level. We therefore adopt Cstat as a conservative empirical threshold for model inclusion throughout this work, including the physically motivated models in Section III.2. This criterion is intended for model selection rather than as a formal line-detection significance test, which would require dedicated Monte Carlo simulations (e.g., Kaastra, 2017).
The phenomenological fits are used to identify statistically significant spectral structure and to guide the construction of the physically motivated models presented in Section III.2. Given the limited spectral resolution and photon statistics of the ACIS-S data, the Gaussian components should be interpreted as representations of blended line complexes rather than formal detections of individual atomic transitions.
@@ III.1 Phenomenological Models
The spectrum shows a pronounced hump between 3-6 keV, along with distinct emission features at both lower and higher energies. The best-fit model (Cstat/d.o.f. = 0.76) requires a myts scattered continuum, its associated mytl fluorescent line component, and an additional Gaussian at keV. Below 3 keV, several Gaussian lines are detected at keV (Mg XI, possibly blended with Mg XII), keV (Si XIII/Si XIV), keV (S XV/S XVI), and keV (S XV). The 7.14 keV feature is discussed further in Section VI.2 as a candidate blueshifted Fe XXVI Ly line. The best-fit line-of-sight absorption is cm-2.
III.2 Physically Motivated Models
To investigate the physical origin of the X-ray emission, we next fit models composed of photoionized and, when necessary, thermal plasma components. The photoionized emission is modeled using Cloudy C.23 (Ferland et al., 1998; Chatzikos et al., 2023), while thermal emission is tested with Apec (Foster et al., 2012). All fits adopt the phenomenological continuum from Section III.1, with the intrinsic absorbing column allowed to vary above the Galactic value.
The Cloudy grids assume a broken power-law spectral energy distribution (SED) of the form , with below 13.6 eV, between 13.6 eV and 0.5 keV, from 0.5-10 keV, and from 10-100 keV, adopting 1.4 solar abundances (Trindade Falcão et al., 2021b, see also Section V.1). The final grids span in steps of 0.25 dex and in 0.1 dex steps. The spectra are exported as XSPEC-style additive tables (Porter et al., 2006) and fitted using Sherpa.
A physical requirement of the photoionization models is that the emitting slab must be geometrically thin at the distance of the gas. Specifically, the slab thickness should satisfy (e.g., Kraemer et al., 2015). This ensures that the cloud does not exceed either the amount of material available at that radius or the maximum emitting surface implied by the morphology. In practice, this condition introduces a coupled constraint between the ionization parameter () and the allowed range of column densities (): for a given , only certain are physically plausible.
Because varies during the fitting process, the upper limit on the column density, , cannot be imposed as a single fixed boundary. Instead, we implement a dynamic constraint in Sherpa that updates the maximum allowed column density at each iteration of the fit. The constraint has the general form
@@ III.2 Physically Motivated Models
The physically motivated models were constructed using the same iterative strategy and statistical criteria adopted for the phenomenological fits in Section III.1. Starting from the minimal photoionized model required to reproduce the dominant soft X-ray emission, additional components (either Cloudy slabs or a thermal Apec plasma) were introduced one at a time and retained only if their inclusion resulted in a statistically meaningful improvement of the fit and a clear reduction of coherent residuals, as described in the following.
For each source, we explored models including up to three photoionized components, as well as combinations of photoionized and thermal plasma components. As in the phenomenological analysis, model complexity was increased only when justified by the data, adopting the same Cstat criterion described in Section III.1 as a conservative empirical threshold for model inclusion. Components that did not meet this criterion, or whose parameters became unconstrained or degenerate with existing components, were not retained.
In FIRST J120041, the data require two distinct photoionized components to reproduce the observed spectrum, whereas the addition of further photoionized or thermal plasma components does not yield a statistically significant improvement nor remove structured residuals. In contrast, the soft X-ray emission of 2MASX J130038 is adequately described by a single photoionized component. In both sources, no thermal plasma component is required by the data, indicating that photoionization dominates the soft X-ray emission. The adopted models therefore provide a sufficient and physically self-consistent description of both spectra.
The resulting best-fit physical models are presented in Figure 2, and their parameter values in Table 3. Here and throughout the paper, parameters labeled as “n.c.” indicate values that are formally unconstrained by the data within the physically allowed parameter space.
@@
The spectrum is best described (Cstat/d.o.f. = 1.73) by one medium-ionization () and medium-column ( cm-2), and one high-ionization () and low-column ( cm-2) photoionized components. No additional continuum is required to reproduce the data.
@@
IV Morphological Properties of the X-ray Emission
IV.1 PSF Modeling
Accurate modeling of the Chandra point spread function (PSF) is essential for distinguishing real extended emission from unresolved nuclear components. For each observation, we simulate the PSF using the Chandra Ray Tracer (ChaRT999https://cxc.cfa.harvard.edu/ciao/PSFs/chart2/index.html) and projected the rays onto the detector plane with MARX101010https://space.mit.edu/cxc/marx/. We generate ensembles of 1,000 PSF realizations centered at the source position and merge them within the same energy bands adopted for the imaging analysis (Section IV.2).
Following the empirical method of Fabbiano et al. (2020), each simulated PSF is normalized to the observed counts within an annulus of 0.2′′-0.5′′ radius. This normalization assumes that the observed emission comprises a central point-like component plus any additional extended structure. By testing multiple normalization radii we ensure robust scaling and minimize the risk of over-subtracting flux from the PSF wings.
IV.2 Broad- and Narrow-Band Imaging